# PFML: Self-Supervised Learning of Time-Series Data Without Representation Collapse

### v2

@@ 7 Conclusion
Table 4: A listing of pros and cons for each of the SSL methods used in the present experiments. For further details on these results, refer to Appendix D.
Limitations
@@
We selected the present set of 11 functionals for their effectiveness across the three data modalities used in the present study, aiming for potential generalizability and a robust starting point to other data domains and downstream tasks. However, carefully selecting the number and type of functionals specifically for different modalities may lead to better results than presented here. Also, we did not include data augmentation in our pre-training processes to save computational time for PFML pre-training, as we wanted to pre-compute the functionals before the model training. As shown in e.g. Chen et al. (2020); Grill et al. (2020); He et al. (2022); Balestriero et al. (2023), data augmentation during pre-training may lead to improved performance on downstream tasks. Nonetheless, performing masking for randomly sampled frames is already a form of data augmentation in itself. Furthermore, other model architectures besides CNN-based encoders or Transformer encoder blocks could also be used, and this may improve PFML pre-training performance. Lastly, we acknowledge that typically SSL pre-training is run with very large minibatch sizes using multiple GPUs, and the results of the present experiments might improve with larger minibatch sizes. However, to promote reproducibility and encourage other researchers to try PFML, we deliberately pre-trained our models using relatively small minibatches so that the pre-training processes could be run on a single GPU with 16 GB of VRAM. As detailed in Appendix E, our method used only a moderate amount of computational resources.
Broader Impacts
@@
Since the main goal of PFML is to make the algorithm straightforwardly applicable to different time-series data domains, our method makes it easier to apply SSL pre-training for time-series data without complex tuning of hyperparameters or the need to profoundly understand the target data domain. As an example, properties of different medical time-series data, such as those obtained with EEG, ECG, or EMG, can be dependent on the clinical environment, the specific measurement equipment and setup, or clinical population being measured (Watson et al., 2019). This limits the applicability of ’universal’ pre-trained models predominant in computer vision and speech technology. In a similar manner, various industrial sensor setups, such as those for system monitoring and predictive maintenance (accelerometers, magnetometers etc.), can result in data unique to a particular environment or machine type. In these cases, the use of PFML pre-training can be practical, since applying modality-specific SSL algorithms or fine-tuning pre-trained models from other data modalities might not generalize well to novel time-series data domains. Hence, PFML may promote the use of machine learning as an assisting tool in e.g. clinical healthcare or other limited-data domains. However, as with all classifiers, machine-learning models trained using PFML might make errors. Incorrect model-based decisions, such as incorrect diagnoses, may be detrimental in some cases. Lastly, any bias, private information, or harmful content in the pre-training data can, in theory, be reflected to the feature representations that are learned by PFML.
8 Acknowledgements
@@
Manu Airaksinen was supported by the Research Council of Finland (grant no. 343498) and the Sigrid Juselius Foundation. Einari Vaaras would like to thank the Finnish Foundation for Technology Promotion for the encouragement grants and the Nokia Foundation for the Nokia Scholarship. The authors would also like to thank the Tampere Center for Scientific Computing for the computational resources used in this study.
Appendix A Proof of Non-Collapsed Feature Representations in PFML Pre-Training
This section provides a more detailed mathematical formulation for the proof that PFML pre-training does not converge to collapsed feature representations.
Let $\mathbf{x}$ be a single- or multi-channel time-series signal, framed into a sequence of short-term frames $\{\mathbf{x}_{0},\mathbf{x}_{1},...\}$ of $N$ samples each, where $\mathbf{x}_{n}=\{x_{t},x_{t+1},...,x_{t+N-1}\}$. We define a set of $m$ functionals, =$\mathcal{F}=\{F_{0},F_{1},...,F_{m-1}\}$, to be computed for each frame $\mathbf{x}_{n}$ to produce a set of computed functionals $\mathbf{f}_{n}$. Here, we refer to functionals as mathematical operations which map a time series of arbitrary length into a single value, such as the mean or variance of the signal. Also, let $\mathbf{z}_{n}$ be the output embeddings of an encoder model given the input $\mathbf{x}_{n}$, and let $\mathbf{y}_{n}$ denote the output predictions of a Transformer-based model given the input $\mathbf{z}_{n}$.
To formalize the relationships between inputs and outputs, let us define the following functions:
  • Let $\mathcal{F}$ be the set of functionals that maps the input frames $\mathbf{x}_{n}$ to the computed functionals $\mathbf{f}_{n}$, i.e., $\mathbf{f}_{n}=\mathcal{F}(\mathbf{x}_{n})=\{F_{0}(\mathbf{x}_{n}),F_{1}(\mathbf{x}_{n}),...,F_{m-1}(\mathbf{x}_{n})\}$.
  • Let $g$ be the function that maps the embeddings $\mathbf{z}_{n}$ to the predictions $\mathbf{y}_{n}$, i.e., $\mathbf{y}_{n}=g(\mathbf{z}_{n})$.
Let us assume the following in PFML pre-training:
  • Assumption 1: There is temporal variability across the frames $\mathbf{x}_{n}$. Formally, let $\sigma^{2}(\mathbf{x}_{n})$ denote the variance of $\mathbf{x}_{n}$ across the frames, and $\sigma^{2}(\mathbf{x}_{n})>0$.
  • Assumption 2: Given Assumption 1, the set of non-trivial functionals $\mathcal{F}$ computed from $\mathbf{x}_{n}$ also contains variance across the frames. Formally, let $\sigma^{2}(\mathbf{f}_{n})$ denote the variance of $\mathbf{f}_{n}$ across the frames, and $\sigma^{2}(\mathbf{f}_{n})>0$.
Under these assumptions, we aim to show that the predictions $\mathbf{y}_{n}$ also contain variance across the frames, i.e., $\sigma^{2}(\mathbf{y}_{n})>0$.
In PFML pre-training, the model learns to predict the computed functionals $\mathbf{f}_{n}$ given the embeddings $\mathbf{z}_{n}$. The prediction loss $L$ is defined as
$L=\frac{1}{N}\sum_{n=1}^{N}\left(\mathbf{f}_{n}-\mathbf{y}_{n}\right)^{2}$ (3)
for the MSE loss and
$L=\frac{1}{N}\sum_{n=1}^{N}\left|\mathbf{f}_{n}-\mathbf{y}_{n}\right|$ (4)
for the L1 loss.
To minimize either the MSE or L1 loss functions (Equations 3 and 4, respectively), the predictions $\mathbf{y}_{n}$ must closely match the computed functionals $\mathbf{f}_{n}$. If $\mathbf{y}_{n}$ were to contain zero variance across the frames, i.e., $\sigma^{2}(\mathbf{y}_{n})=0$, while $\mathbf{f}_{n}$ contains variance, i.e., $\sigma^{2}(\mathbf{f}_{n})>0$, the prediction loss $L$ would be high. This is because the constant predictions $\mathbf{y}_{n}$ would not be able to capture the temporal variability in $\mathbf{f}_{n}$.
Therefore, to achieve low prediction loss values, the predictions $\mathbf{y}_{n}$ must also contain variance across the frames, i.e., $\sigma^{2}(\mathbf{y}_{n})>0$. Consequently, PFML pre-training does not converge to collapsed feature representations, as long as Assumptions 1 and 2 hold true. $\square$
Given that real-world time-series data generally shows temporal variability, and computed functionals derived from such data are expected to reflect this variability, Assumptions 1 and 2 are valid for most real-world datasets.
Appendix B Pre-Training and Fine-Tuning Hyperparameters
This section provides details on the pre-training (Table 5) and fine-tuning (Table 6) hyperparameters of the present experiments.
Table 5: The pre-training hyperparameters for PFML, data2vec, and MAE pre-training for each data modality (IMU, speech, and EEG data).
Table 6: The fine-tuning hyperparameters for PFML, data2vec, and MAE pre-trained models for each data modality (IMU, speech, and EEG data).
Appendix C Results on Additional Hyperparameter Experiments
This section provides the result tables for the additional hyperparameter experiments in Section 6. Table 7 presents the fine-tuning results for the comparison between masking either the model inputs or embeddings during PFML pre-training. Tables 8, 9, and 10 show the fine-tuning results for different configurations of masking probabilities ($p_{\text{m}}$) and mask lengths ($l_{\text{m}}$) for multi-sensor IMU, speech, and EEG data, respectively. Table 11 shows the fine-tuning results for discarding some of the 11 functionals in PFML pre-training for multi-sensor IMU data. Finally, Table 12 presents the fine-tuning results for different mask types for PFML pre-training for multi-sensor IMU data.
Table 7: The fine-tuning results for the comparison between masking either inputs or embeddings for PFML pre-training.
Table 8: The fine-tuning results for different PFML pre-training masking hyperparameter configurations for multi-sensor IMU data. The results are shown both with a fixed mask length and varying masking probability (left), and vice versa (right).
Table 9: The fine-tuning results for different PFML pre-training masking hyperparameter configurations for speech data. The results are shown both with a fixed mask length and varying masking probability (left), and vice versa (right).
Table 10: The fine-tuning results for different PFML pre-training masking hyperparameter configurations for EEG data. The results are shown both with a fixed mask length and varying masking probability (left), and vice versa (right).
Table 11: The fine-tuning results for discarding some of the functionals in PFML pre-training for IMU data.
Table 12: The fine-tuning results for different mask types for PFML pre-training for IMU data.
Appendix D Further Details on the Pros and Cons of Each SSL Method
This section provides further details on the results presented in Table 4. Table 13 presents the numerical values behind the results of Table 4.
Table 13: A listing of pros and cons for each of the SSL methods used in the present experiments. Up arrow ($\uparrow$) indicates that a higher score is better, whereas a down arrow ($\downarrow$) indicates that a lower score is better.
The values for the columns in Table 13 have been defined as follows:
  • End-use performance: These results are the average of the UAF1 (%) and UAR (%) values in Table 1 across all five classification tasks.
  • Robust to representation collapse: This is a “Yes/No” decision based on Table 3.
  • Expected pre-training time: The expected pre-training time $t_{\text{pt}}(P_{\text{m}})$ for pre-training method $P_{\text{m}}$ is computed as
    $t_{\text{pt}}(P_{\text{m}})=\left(avg\_num\_collapses+1\right)\cdot avg\_training\_time\ ,$ (5)
    where $avg\_num\_collapses$ is the average number of representation collapses across all three data modalities (Table 3) and $avg\_training\_time$ is the average pre-training time (in hours) across all three data modalities. Note that for MAE, we considered the average number of representation collapses to be zero, since there was only a single representation collapse which was most probably due to bad luck in model weight initialization. Also note that model training time can vary significantly depending on the amount and type of training data, the software implementation, and the computational hardware used. However, when taking the expected number of representation collapses during pre-training into account, the SSL methods used in the present experiments can be categorized into three distinct groups.
  • Flexibility of neural network architecture: This is a “Yes/No” decision based on whether the SSL algorithm has a free choice of the encoder and classifier architecture. In TS2Vec, the pre-training algorithm is built into the pre-defined CNN encoder architecture, and therefore it is not possible select the architecture of the encoder freely.
  • Linear separability of SSL features: These results are the average of the UAF1 (%) and UAR (%) values in Table 2 across all five classification tasks.
  • Number of hyperparameters: These values are a simple count of the number of different SSL algorithm-specific hyperparameters that affect the performance of the SSL algorithm. These algorithm-specific hyperparameters for each SSL algorithm are:
    • TS2Vec: 1) Minibatch size, 2) Number of negative samples in minibatch, 3) Data augmentation method
    • MAE: 1) Masking probability, 2) Mask length
    • data2vec: 1) Masking probability, 2) Mask length, 3) Student network learning rate, 4) Teacher network, initial weight update rate, 5) Teacher network, final weight update rate, 6) Teacher network, number of weight update rate transitions, 7) Training targets
    • PFML: 1) Masking probability, 2) Mask length, 3) Functionals
Appendix E Additional Information on Computational Resources
All computations were run on a computing cluster operating on the SLURM environment. For both model pre-training and fine-tuning, we used an NVIDIA Tesla V100 GPU with 16 GB of VRAM, four CPU cores, and 16 GB of RAM. Table 14 shows the pre-training durations for each data modality (multi-sensor IMU data, speech data, and EEG data). Note that due to RAM constraints, samples of each minibatch were separately loaded from a disk to RAM. The possibility to load the entire training dataset into RAM would speed up the pre-training process substantially.
Table 14: The PFML pre-training durations for each data modality (IMU, speech, and EEG data).
The research project required more computations than what is reported in the present paper: For each data modality and pre-training algorithm, preliminary experiments were conducted in order to find suitable hyperparameters for both the pre-training and fine-tuning processes (Tables 5 and 6, respectively).
Appendix F Additional TS2Vec Results
This section provides all TS2Vec-related results of the present experiments. Table 15 shows a summary of all of the results regarding TS2Vec for IMU, speech, and EEG data, further described in Section 4.5.
Table 15: All TS2Vec-related results across the three different data modalities (IMU, speech, and EEG data). Starting from the top, the table includes results for model fine-tuning with a randomly initialized model consisting of the TS2Vec encoder and Transformer classifier, followed by results for the same model with the encoder pre-trained using TS2Vec. The rest of the results are without the Transformer: First, linear evaluation results for the TS2Vec encoder without pre-training, followed by linear evaluation results using a TS2Vec pre-trained encoder. Finally, following the procedure used in the original TS2Vec paper (Yue et al., 2022), results are provided using an SVM classifier with an RBF kernel for the SSL features obtained from the pre-trained TS2Vec encoder.

### v3

@@ V Results
TABLE IV: A listing of pros and cons for each of the SSL methods used in the present experiments. For further details on these results, refer to Appendix D in the supplementary material.
The results of representation collapse experiments are shown in Table III. As can be seen from the results, it is very common for representation collapse to occur with data2vec across all data modalities. On the contrary, the results indicate that PFML, MAE, and TS2Vec do not suffer from representation collapse: PFML and TS2Vec did not experience representation collapses at all, and MAE had a representation collapse only once. Furthermore, we attribute this single representation collapse of MAE to bad luck in model weight initialization, as in this particular case the model loss started diverging from the beginning of the pre-training process. The results showcase that methods like PFML, MAE, and TS2Vec, whose training targets inherently contain variance, are less prone to representation collapse compared to methods like data2vec that learn their own prediction targets.
[…]
VI Additional Hyperparameter Experiments
In order to demonstrate that it is more beneficial during pre-training to mask the latent features instead of masking the input directly, we ran PFML pre-training for all three datasets twice: either by masking the inputs or by masking the embeddings. Subsequently, we fine-tuned our models for all five classification tasks, and the results are shown in Table 7 of Appendix C in the supplementary material. As can be observed from the results, it is more beneficial for downstream tasks if we alleviate the complexity of the pre-training task for the encoder by masking the embeddings instead of masking the inputs. The only exception was with EEG data, where it did not make a difference whether inputs or embeddings were masked.
For each data modality, we also experimented with different configurations of masking probability $p_{\text{m}}$ and the length of the masks $l_{\text{m}}$. We ran PFML pre-training using different configurations of $p_{\text{m}}$ and $l_{\text{m}}$, and then we fine-tuned the pre-trained models. For IMU and speech data, we only experimented with one classification task each, namely classification of movement from IMU data and classification of valence from speech data. The results for different configurations of $p_{\text{m}}$ and $l_{\text{m}}$ for IMU, speech, and EEG data are shown in Appendix C in the supplementary material in Tables 8, 9, and 10, respectively. For IMU data, the differences between different masking strategies are rather small, whereas for speech and EEG data the selection of masking hyperparameters has a notable effect on fine-tuning performance.
We also experimented with the effect of discarding some of the functionals in PFML pre-training for IMU data. After pre-training, we fine-tuned our model for movement classification, and the results are presented in Table 11 of Appendix C in the supplementary material. The results indicate that using the full set of 11 functionals during PFML pre-training provides the best outcome. As the number of discarded functionals increases, the prediction task becomes simpler and the training targets are able to capture less information of the input signal frames, leading to worse fine-tuning performance.
Finally, we tested different mask types for PFML pre-training using IMU data. We either replaced the masked embeddings with a fixed vector of zeros, ones, random Gaussian noise (as in e.g. Baevski et al. [11]), or a learnable mask token (as in e.g. Baevski et al. [8]). After PFML pre-training using the four different mask types, we fine-tuned the pre-trained models for movement classification. Table 12 of Appendix C in the supplementary material presents the comparison results for different mask types. As can be observed, the choice between a mask of ones or random Gaussian noise does not have a notable impact on the performance. However, using a learnable mask token yielded slightly worse results than a vector of ones or random Gaussian noise, and a vector of zeros yielded the worst results. We observed that either using a vector of ones, random Gaussian noise, or learnable mask tokens for masking the embeddings promoted embedding variance, whereas using a vector of zeros provided a smaller level of variance for the embedding representations during pre-training. This lower level of variance for embeddings might potentially hinder the fine-tuning process, resulting into a lower performance in downstream tasks.
VII Conclusion
In this paper, we presented PFML, a novel SSL algorithm for time-series data that avoids the common SSL issue of representation collapse. PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. We demonstrated the effectiveness of PFML using five different classification tasks across three different data modalities: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. Our results show that PFML is superior to both a conceptually similar SSL method, MAE, and a contrastive learning-based SSL method, TS2Vec. Our results also show that PFML is on par with the current state-of-the-art data modality agnostic SSL method, data2vec, while being conceptually simpler and without suffering from representation collapse. The pros and cons for each of the SSL methods used in the present study are shown in Table IV. The fact that PFML matches the performance of data2vec while also avoiding the issue of representation collapse renders PFML more straightforward to apply to new time-series data domains, such as in the case of clinical time-series data. The present work may also be extended to other domains than time-series data, such as images where functionals could be computed of, e.g., image patches.
VII-A Limitations
@@
We selected the present set of 11 functionals for their effectiveness across the three data modalities used in the present study, aiming for potential generalizability and a robust starting point to other data domains and downstream tasks. However, carefully selecting the number and type of functionals specifically for different modalities may lead to better results than presented here. Also, we did not include data augmentation in our pre-training processes to save computational time for PFML pre-training, as we wanted to pre-compute the functionals before the model training. As shown in e.g. [6, 26, 28, 1], data augmentation during pre-training may lead to improved performance on downstream tasks. Nonetheless, performing masking for randomly sampled frames is already a form of data augmentation in itself. Furthermore, other model architectures besides CNN-based encoders or Transformer encoder blocks could also be used, and this may improve PFML pre-training performance. Lastly, we acknowledge that typically SSL pre-training is run with very large minibatch sizes using multiple GPUs, and the results of the present experiments might improve with larger minibatch sizes. However, to promote reproducibility and encourage other researchers to try PFML, we deliberately pre-trained our models using relatively small minibatches so that the pre-training processes could be run on a single GPU with 16 GB of VRAM. As detailed in Appendix E in the supplementary material, our method used only a moderate amount of computational resources.
VII-B Broader Impacts
@@
Since the main goal of PFML is to make the algorithm straightforwardly applicable to different time-series data domains, our method makes it easier to apply SSL pre-training for time-series data without complex tuning of hyperparameters or the need to profoundly understand the target data domain. As an example, properties of different medical time-series data, such as those obtained with EEG, ECG, or EMG, can be dependent on the clinical environment, the specific measurement equipment and setup, or clinical population being measured [60]. This limits the applicability of “universal” pre-trained models predominant in computer vision and speech technology. In a similar manner, various industrial sensor setups, such as those for system monitoring and predictive maintenance (accelerometers, magnetometers etc.), can result in data unique to a particular environment or machine type. In these cases, the use of PFML pre-training can be practical, since applying modality-specific SSL algorithms or fine-tuning pre-trained models from other data modalities might not generalize well to novel time-series data domains. Hence, PFML may promote the use of machine learning as an assisting tool in e.g. clinical healthcare or other limited-data domains. However, as with all classifiers, machine-learning models trained using PFML might make errors. Incorrect model-based decisions, such as incorrect diagnoses, may be detrimental in some cases. Lastly, any bias, private information, or harmful content in the pre-training data can, in theory, be reflected to the feature representations that are learned by PFML.
Acknowledgment
@@
The authors would like to thank Tampere Center for Scientific Computing for the computational resources used in this study. The author Einari Vaaras would like to thank the Finnish Foundation for Technology Promotion for the encouragement grants and the Nokia Foundation for the Nokia Scholarship.